OpenStudy (anonymous):

[______] find the perimeter length = 8x^4+5x^3 width = 2x^3-x

OpenStudy (anonymous):

2(8x^4+7x^3-x) i.e 16x^4+14x^3-2x

OpenStudy (shinalcantara):

Perimeter = 2L + 2W hence, \[Perimeter = 2(8x ^{4} + 5x ^{3}) + 2(2x ^{3} - x)\] \[Perimeter = 16x ^{4} + 10x ^{3} + 4x ^{3} - 2x\] \[Perimeter = 16x ^{4} + 14x ^{3} - 2x\]